APPENDIX B: Complete Sample Technical TEX Document

This appendix will eventually consist of TEX output for “A Computationally Fast Implementation of Fisher’s Test of Difference-Score Population Symmetry About Zero.”

That manuscript has been submitted for publication in a statistical journal, however unanticipated delays in the publication of the article prevent me from including a copy of it as Appendix B at this time.

As soon as that article goes to press, a copy of the TEX output and the commands used to generate it will be made available so that TEX users can have a technical article (complete with equations) to use as an example for their own technical articles.

I regret any inconvenience this unavoidable delay may cause.